OpenStudy (anonymous):

solve for x: (x+5)(x+2)/(x+1)= -1

OpenStudy (anonymous):

\[ \{-4+\sqrt {5},-4-\sqrt {5}\} \]

OpenStudy (anonymous):

do you want the steps?

OpenStudy (anonymous):

\[x^2+7x+10=-x-1\]\[x^2+8x+11=0\]use quadratic formula or (as I will do) complete the square to solve\[x^2+8x+16=-11+16\]\[(x+4)^2=5\]\[x=-4 \pm \sqrt{5}\]